# An improved lower bound for Erdős–Szekeres products

### v1

@@ Abstract.
This gives the first improvement of the classical lower bound for the Erdős–Szekeres problem in the general case since 1959.
@@ 1. Introduction
Among the many problems raised by Erdős and Szekeres, the most central one is the following, as highlighted in [10, Problem 1] and also listed as Problem #256 on Bloom’s Erdős Problems website [15]:
@@
On the lower bound side, the classical result $f(n)\geq\sqrt{2n}$ goes back to Erdős and Szekeres [8]. Although their original proof contained a gap, this was later repaired in [3] (see also Section 2 for details). In 2021, Lubinsky et al. [3, Theorem 2.1] obtained new lower bounds using the Poisson integral of $\log|P_{n}|$, which are effective when the $\{s_{j}\}$ do not grow too quickly. Another result in the same paper, [3, Theorem 2.2], proved using Kellogg’s extension of the Hausdorff–Young inequalities for Fourier coefficients, gives good bounds when the $\{s_{j}\}$ grow rapidly or are well separated. Nevertheless, as emphasized in [10, p. 300], there has been no improvement on the lower bound $f(n)\geq\sqrt{2n}$ in general.
Our main result gives a genuine improvement of the classical lower bound in general:
@@ 2. The classical lower bound
following the exposition in [3, Section 3]. This will make it clear how our new approach strengthens the existing argument.
@@ 2.1. The original Erdős–Szekeres idea
Erdős [8] and Szekeres considered the identity
@@ 2.3. A corrected proof
A rigorous proof was later given in [3, Section 3]. Since $(1-z)^{n}\mid P_{n}(z)$, we have
@@
for every polynomial $S(x)$ of degree at most n1n-1. As $P_{n}$ is not the zero polynomial, this forces at least $n$ coefficients $a_{k}$ to be nonzero. [^1] In fact, with a more refined argument one can show that $P_{n}$ has at least $n+1$ nonzero coefficients. Hence $\sum_{k=0}^{N}|a_{k}|^{2}\geq n$. Finally, by a result of O’Hara and Rodriguez [14, Corollary 1], we have
@@
This is the classical lower bound, rigorously justified in [3]. Interestingly, in [3] the authors further remarked:
@@
Our present work may be viewed as providing precisely such a fix, thereby confirming their expectation. Indeed, the proof of Theorem 1.2 also makes use of the O’Hara–Rodriguez bound. The key new ingredient, however, is that in Theorem 3.2 we prove the stronger estimate
@@ 3. Proof of the Main Result
For the polynomials of the form (1.1), Maltby [11, 12] and Borwein [5, p. 104] obtained some improvements for specific values of $n$, focusing on the $\ell^{1}$ norm of the coefficients and employing algorithms to search for optimal choices of $\{s_{j}\}$. It is worth noting that in [12, p. 1323] Maltby pointed out the inequality $\|P_{n}\|_{1}\geq 2n$, referring to [6] for a proof. Similarly, in Borwein [5, p. 104], the inequality $\|P_{n}\|_{1}\geq 2n$ was also stated, arguing that it follows directly from the Prouhet–Tarry–Escott (PTE) problem. However, these statements are limited in scope: just as in the gap in Erdős and Szekeres [8], both [6] and [5] applied the Prouhet–Tarry–Escott problem, which is only valid for the special form
@@
Moreover, by [14, Corollary 1], one knows that

### v2

@@ Abstract.
This gives the first improvement of the classical lower bound for the Erdős–Szekeres problem in the general case since 1959. In particular, our result confirms a remark of Billsborough et al. [4], who observed that if the original Erdős–Szekeres proof could be fixed, the O’Hara–Rodriguez bound would yield exactly this inequality.
@@ 1. Introduction
Among the many problems raised by Erdős and Szekeres, the most central one is the following, as highlighted in [13, Problem 1] and also listed as Problem #256 on Bloom’s Erdős Problems website [6]:
@@
On the lower bound side, the classical result $f(n)\geq\sqrt{2n}$ goes back to Erdős and Szekeres [11]. Although their original proof contained a gap, this was later repaired in [4] (see also Section 2 for details). There have also been several results that treat the sequence $\{s_{j}\}$ under additional restrictions; see [13, pp. 300–302] for a detailed account, of which we give only a brief summary here. Borwein [7] obtained a new lower bound in the case when none of the $\{s_{j}\}$ are divisible by a given prime $p$. Bell, Borwein, and Richmond [2] obtained both asymptotic upper and lower bounds. Bourgain and Chang [9], using probabilistic methods and careful estimation of trigonometric sums, established new results for “dense” sets $S=\{s_{1}<\cdots<s_{n}\}$. In 2021, Billsborough et al. [4, Theorem 2.1] obtained new lower bounds using the Poisson integral of $\log|P_{n}|$, which are effective when the $\{s_{j}\}$ do not grow too quickly. Another result in the same paper, [4, Theorem 2.2], proved using Kellogg’s extension of the Hausdorff–Young inequalities for Fourier coefficients, gives good bounds when the $\{s_{j}\}$ grow rapidly or are well separated. Nevertheless, as emphasized in [13, p. 300] (see also [9, p. 253] or [4]), there has been no improvement on the lower bound $f(n)\geq\sqrt{2n}$ in general.
Our main result is the following theorem:
@@ 2. The classical lower bound
following the exposition in [4, Section 3]. This will make it clear how our new approach strengthens the existing argument.
@@ 2.1. The original Erdős–Szekeres idea
Erdős and Szekeres [11] considered the identity
@@ 2.3. A corrected proof
A rigorous proof was later given in [4, Section 3]. Since $(1-z)^{n}\mid P_{n}(z)$, we have
@@
for every polynomial $S(x)$ of degree at most n1n-1. As $P_{n}$ is not the zero polynomial, this forces at least $n$ coefficients $a_{k}$ to be nonzero. [^1] In fact, with a more refined argument one can show that $P_{n}$ has at least $n+1$ nonzero coefficients. Hence $\sum_{k=0}^{N}|a_{k}|^{2}\geq n$. Finally, by a result of O’Hara and Rodriguez [17, Corollary 1], we have
@@
This is the classical lower bound, rigorously justified in [4]. Interestingly, in [4] (see also [13, p. 298]) the authors further remarked:
@@
Our present work may be viewed as providing precisely such a fix, thereby confirming their remark. Indeed, the proof of Theorem 1.2 also makes use of the O’Hara–Rodriguez bound. The key new ingredient, however, is that in Theorem 3.2 we prove the stronger estimate
@@ 3. Proof of the Main Result
For the polynomials of the form (1.1), Maltby [14, 15] obtained some improvements for specific values of $n$, focusing on the $\ell^{1}$ norm of the coefficients and employing algorithms to search for optimal choices of $\{s_{j}\}$. It is worth noting that in [15, p. 1323] Maltby pointed out the inequality $\|P_{n}\|_{1}\geq 2n$, referring to [8] for a proof. However, this statement is limited in scope: just as in the gap in Erdős and Szekeres [11], [8, Section 4.2] applied the Prouhet–Tarry–Escott problem, which is only valid for the special form (in [8, Section 4.2] also referred to as a pure product)
@@
Moreover, by [17, Corollary 1], one knows that